# Multicritical quantum sensors driven by symmetry-breaking

### v1

@@ Appendix B Derivation of $\mathcal{F}_{ab}$ for ABC
For the Kitaev model the ground state is of the form [48] :
@@
where $k$ stands for the momentum states, $k=(2i-1)\pi/L$, for $i=1,..,L/2$. For brevity, we mark $k$ via $i$ in the product above and $\gamma_{k}^{\dagger}=u_{k}+v_{k}c_{k}^{\dagger}c_{-k}^{\dagger}.$ Furthermore,
@@
$\Rightarrow\langle\partial_{a}\Psi_{G}|\Psi_{G}\rangle$ $=\sum_{k=1}^{L/2}\partial_{a}u_{k}^{*}u_{k}+\partial_{a}v_{k}^{*}v_{k}.$ (14)
@@
$+4\text{Re}\left[\left(\sum_{k=1}^{L/2}\partial_{a}u_{k}^{*}u_{k}+\partial_{a}v_{k}^{*}v_{k}\right)\left(\sum_{k=1}^{L/2}u_{k}^{*}\partial_{b}u_{k}+v_{k}^{*}\partial_{b}v_{k}\right)\right].$ (15)
We note that $\sum_{k=1}^{L/2}\partial_{a}u_{k}^{*}u_{k}+\partial_{a}v_{k}^{*}v_{k}=0$ using $|u_{k}|^{2}+|v_{k}|^{2}=1$. Thus, we get:
@@
$\partial_{\Delta}u_{k}$ $=\frac{-y_{k}^{2}z_{k}}{2\sqrt{2}\Delta e^{5/2}_{k}\sqrt{\mathcal{E}_{k}+z_{k}}},$
$\partial_{\Delta}v_{k}$ $=\frac{iy_{k}z_{k}(\mathcal{E}_{k}+z_{k})}{2\sqrt{2}\Delta e^{5/2}_{k}\sqrt{\mathcal{E}_{k}+z_{k}}}.$ (17)
@@
where $k=(2i-1)\pi/L$ and sum is over the values $i={1,2,..,L/2}$.
Figure B.1: Comparison between analytic expression (18) of the QFIM and their corresponding values obtained numerically: The elements of QFIM are presented with change in $\mu$ or $\Delta$, i.e., (a) $\mathcal{F}_{\mu\mu}$ against $\mu$, (b) $\mathcal{F}_{\Delta\Delta}$ against $\Delta$, (c)$\mathcal{F}_{\mu\Delta}$ against $\mu$ and (d) $\mathcal{F}_{\mu\Delta}$ against $\Delta$. This is a representative case of system-size $L=1000$, where $\Delta=0.001$ in (a) and (c), while $\mu=2.0$ in (b) and (d). The blue lines are obtained from the Eq.(18), while the orange circular markers are obtained numerically for the Kitaev chain with anti-periodic boundary condition. The y-axes of (a),(b) and (d) are in log-scale while the y-axis of (c) is in linear-scale. The x-axes of all four are in linear-scale.
Figure B.2: Comparison between (18) and (28): The analytic form of the elements of the Quantum Fisher Information matrix: (a) $\mathcal{F}_{\mu\mu}$, (b) $\mathcal{F}_{\Delta\Delta}$, (c) $\mathcal{F}_{\mu\Delta}$ and (d) $\mathcal{G}$ are presented with respect to $\Delta$ which is in the range $\Delta\in[10^{-8},10^{-6}]$, keeping a fixed $\mu=2$. We numerically confirm that these quantities (18) and its approximations (28) in the small $\Delta$ regime matches. The markers indicate the approximated functions (28) and the lines correspond to (18). These are plotted for system sizes $L=400$ (circular markers, dot-dashed line) , $700$ (square markers, solid line) and $1000$ (triangular markers, dotted line). The markers and line-styles are uniform across all the four sub-figures. The y-axes are in log-scale in (a),(b) and (d) and in linear-scale in (c). The x-axis is in linear-scale and the labels are multiplied by $10^{7}$ for convenience.
@@
$\zeta(2)=\sum_{n=1}^{\infty}\frac{1}{n^{2}}=\sum_{k}=\frac{\pi^{2}}{6},$
@@
For very large $L$ and consequently large $m$, we can approximate the sum in bracket of (23) with (24). We thus get,
@@
Figure B.3: Finite-size scaling of elements QFIM and $\mathcal{G}$: We evaluate $\mathcal{F}_{\mu\mu}$ (circular markers), $\mathcal{F}_{\Delta\Delta}$ (square markers), $-\mathcal{F}_{\mu\Delta}$ (triangular markers) and $\mathcal{G}$ (star markers) at the point $\Delta=10^{-10}$ and $\mu=2.0$ using Eq.(18) for system-sizes from $L=1000$ to $10000$. The scaling exponents match with the ones obtained in Eq.(28). The lines represent the best fits.
Figure B.4: Elements of the Quantum Fisher Information matrix: The QFIM elements (a) $\mathcal{F}_{\mu\mu}$ and (b) $\mathcal{F}_{\Delta\Delta}$ are presented with respect to the tuning parameters $\mu$ and $\Delta$ in the vicinity of the critical points of either axes. The probe considered is the ground state of the Kitaev model on a one-dimensional lattice with system size of $L=1000$. Note that in (b) the Fisher information is in the range of $10^{3}$. In both figures, the elements of QFIM are presented in log-scale.
Figure B.5: Change in the average particle number $\langle N\rangle$ with change in $\Delta$: The average particle number of the probe state (11) is numerically obtained and is presented with a fixed $\mu=2$ throughout. This a representative case for three system sizes : (a) L = 400 (dot-dashed line), (b) L = 600 (solid line) and (c) L = 800 (dotted line).
Thus, we have $-\mathcal{F}_{\mu\Delta}\sim L^{4}$ for $\Delta\rightarrow 0$ and $\mu\rightarrow 2$. Now, for the case of $\mathcal{G}$, it can be shown that
@@
$\mathcal{F}_{\mu\mu}$ $\approx\frac{\Delta^{2}}{\pi^{6}}L^{6},$
$\mathcal{F}_{\Delta\Delta}$ $\approx\frac{1}{8}L^{2},$
$\mathcal{F}_{\mu\Delta}$ $\approx-\frac{\Delta}{\pi^{4}}L^{4},$
$\mathcal{G}\approx$ $\frac{(\pi^{2}-8)\Delta^{2}}{\pi^{8}}L^{6}.$ (28)
@@
In Fig. B.4 we present (a) $\mathcal{F}_{\mu\mu}$ and $(b)\mathcal{F}_{\Delta\Delta}$ for system-size $L=1000$. We observe a similar nature for other system-sizes as well. Figure B.4 quite clearly shows the transition from topological to trivial phase at $\mu=2.0$. We see in Fig. B.4(a) that the system achieves its maximum $\mathcal{F}_{\mu\mu}$ at the point $\mu=2.0$. The quantity $\mathcal{F}_{\mu\mu}$ gradually decreases as we move further into the trivial phase. Due to the apparent fluctuations of $\mathcal{F}_{\mu\mu}$ in the topological phase of the finite-sized systems, particularly in smaller systems, numerical calculations were not pushed in the smaller $\Delta$ regime. We, however, performed the scaling analysis in the trivial phase away from the critical line. In Fig. B.4(b), we present the component $\mathcal{F}_{\Delta\Delta}$ of QFIM. The figure clearly shows that this quantity increases as we move nearer to $\Delta=0$ line, across which the system the topological invariant undergoes a discrete jump, i.e., a flip in sign of the winding number, given $|\mu|\leq 2$.
@@
In Fig. B.5, we present the average particle number $\langle N\rangle$ of the probe state, i.e., the ground state of the Kitaev model in one dimension for $\mu=2$. As seen from the Hamiltonian of the Kitaev model, the $p$-wave superconductivity term with coupling $\Delta$ makes the system particle-number non-conserving. This means that when $\Delta$ is set to a non-zero value, the ground state wave-function becomes a superposition of various particle-number sector. When $\Delta$ is set to zero, we note two things: the states with fixed momentum become eigenstates of the Hamiltonian and the value of $\mu$ determines the fixed particle-number sector to which the ground state belongs to. We note that at $\mu=0$ limit, half of the single-particle states have negative eigenenergies. The many-body ground state of the model is constructed by filling the negative energy states via fermions. Since there is a negative sign with $\mu$ in the Hamiltonian, more single-particle eigenstates with negative energies becomes available with increasing $\mu$. Thus, occupation number of the ground state increases gradually with increasing $\mu$. On introducing a non-zero $\Delta$, the ground state no longer has a fixed number of particles, rather it is a superposition of many fixed particle-number states. We observe that at very small values of $\Delta$, the ground state has the fully-filled state with almost unit probability because of a small $\Delta$ and a very large $\mu$. As the value of $\Delta$ increases, it becomes energetically more costly to fill up particles than that for smaller $\Delta$. Due to this, the average particle-number decreases with increase of $\Delta$. Hence, physically it turns out that at the gapless multicritical point ($\mu=2$ and $\Delta=0$), the system is in a trivial band-insulating phase, as all available energy states are filled-up by the fermions. A sufficiently small value of the symmetry-breaking term $\Delta$, i.e., $\Delta^{*}$, ensures the system’s transition to a critical $p$-wave superconducting phase. This, in turn, facilitates the quantum enhanced sensing despite the fact that adiabatic variation of the parameter, $\Delta$, follows the gapless line throughout.

### v2

@@ Appendix A Derivation of $\mathcal{F}_{ab}$
For the Kitaev model the ground state is of the form:
@@
where $k$ stands for the momentum states, $k=(2i-1)\pi/L$, for $i=1,..,L/2$. Here we do the calculations considering the anti-periodic boundary conditions (ABC), but these results holds for periodic boundary conditions (PBC) as well. For brevity, we mark $k$ via $i$ in the product above and $\gamma_{k}^{\dagger}=u_{k}+v_{k}c_{k}^{\dagger}c_{-k}^{\dagger}.$ Furthermore,
@@
$\Rightarrow\langle\partial_{a}\Psi_{G}|\Psi_{G}\rangle$ $=\sum_{k=1}^{L/2}u_{k}\partial_{a}u_{k}^{*}+v_{k}\partial_{a}v_{k}^{*}.$ (19)
@@
$+4\text{Re}\left[\left(\sum_{k=1}^{L/2}u_{k}\partial_{a}u_{k}^{*}+v_{k}\partial_{a}v_{k}^{*}\right)\left(\sum_{k=1}^{L/2}u_{k}^{*}\partial_{b}u_{k}+v_{k}^{*}\partial_{b}v_{k}\right)\right].$ (20)
We note that $\sum_{k=1}^{L/2}u_{k}\partial_{a}u_{k}^{*}+v_{k}\partial_{a}v_{k}^{*}=0$ using $|u_{k}|^{2}+|v_{k}|^{2}=1$ and Eq.  (A). Thus, we get:
@@
$\partial_{\Delta}u_{k}$ $=\frac{-y_{k}^{2}z_{k}}{2\sqrt{2}\Delta\mathcal{E}_{k}^{5/2}\sqrt{\mathcal{E}_{k}+z_{k}}},$
$\partial_{\Delta}v_{k}$ $=\frac{iy_{k}z_{k}(\mathcal{E}_{k}+z_{k})}{2\sqrt{2}\Delta\mathcal{E}_{k}^{5/2}\sqrt{\mathcal{E}_{k}+z_{k}}}.$ (22)
@@
where $k=(2i-1)\pi/L$ and sum is over the values i=1,2,..,L/2i={1,2,..,L/2}. We corroborate the expressions obtained in Eq. (23) by comparing them with numerically obtained values of the elements of the QFI in Fig. 7.
[…]
Figure 8: Comparison between Eq. (23) and Eq. (33): The analytic form of the elements of the Quantum Fisher Information matrix: (a) $\mathcal{F}_{\mu\mu}$, (b) $\mathcal{F}_{\Delta\Delta}$, (c) $\mathcal{F}_{\mu\Delta}$ and (d) $\mathcal{G}$ are presented with respect to $\Delta$ which is in the range $\Delta\in[10^{-8},10^{-6}]$, keeping a fixed $\mu=2$. We numerically confirm that these quantities (from Eq. (23)) and its approximations (from Eq. (33)) in the small $\Delta$ regime matches. The markers indicate the approximated functions as written in Eq. (33) and the lines correspond to Eq. (23). These are plotted for system sizes $L=400$ (circular markers, dot-dashed line) , $700$ (square markers, solid line) and $1000$ (triangular markers, dotted line). The markers and line-styles are uniform across all the four sub-figures. The y-axes are in log-scale in (a),(b) and (d) and in linear-scale in (c). The x-axis is in linear-scale and the labels are multiplied by $10^{7}$ for convenience.
@@
Figure 9: Finite-size scaling of elements QFIM and $\mathcal{G}$: We evaluate $\mathcal{F}_{\mu\mu}$ (circular markers), $\mathcal{F}_{\Delta\Delta}$ (square markers), $\mathcal{F}_{\mu\Delta}$ (triangular markers) and $\mathcal{G}$ (star markers) at the point $\Delta=10^{-10}$ and $\mu=2.0$ using Eq. (23) for system-sizes from $L=1000$ to $10000$. The scaling exponents match with the ones obtained in Eq. (33). The lines represent the best fits.
Figure 10: Elements of the Quantum Fisher Information matrix: The QFIM elements (a) $\mathcal{F}_{\mu\mu}$ and (b) $\mathcal{F}_{\Delta\Delta}$ are presented with respect to the tuning parameters $\mu$ and $\Delta$ in the vicinity of the critical points of either axes. The probe considered is the ground state of the Kitaev model on a one-dimensional lattice with system size of $L=1000$. Note that in (b) the Fisher information is in the range of $10^{3}$. In both figures, the elements of QFIM are presented in log-scale.
Figure 11: Validity of the small $\Delta$ expansion: We observe that the system-size scaling of $\mathcal{F}_{\mu\mu}$ undergoes a gradual transition from $L^{6}$ to $L^{2}$ as system-size $L$ increases. We define a specific system-size $L_{c}$ (represented by black dots and dotted-vertical lines), at which value the actual $\mathcal{F}_{\mu\mu}$ deviates from the corresponding $L^{6}$ fit by $1\%$. We realize that $L_{c}$ increases as $\Delta$ decreases. This validates our assumption of considering $\Delta\rightarrow 0$ limit as we can always find a finite $L_{c}$ for a corresponding non-zero $\Delta$. In the inset, we fit the $L_{c}$ with $\Delta$ and find the fit : $L_{c}\sim\Delta^{-1}$.
Validity of $\mathcal{F}_{\mu\mu}\sim L^{6}$ : In Fig. 11, we observe that for a specific value of $\Delta$, there exist a $L$ value, up to which we observe $\mathcal{F}_{\mu\mu}\sim L^{6}$ and beyond which $\mathcal{F}_{\mu\mu}\sim L^{2}$. In the Fig. 11, we present this for different values of $\Delta$. We define a specific system-size $L_{c}$, which is the system-size at which the actual value of $\mathcal{F}_{\mu\mu}$ deviates by $1\%$ from the extrapolated value it would have had with $\mathcal{F}_{\mu\mu}\sim L^{6}$ fit. The black dotted lines on the $L$-axis, corresponds to the different critical lengths $L_{c}$-s obtained for various values of $\Delta$-s. In the inset we observe that they fit as $L_{c}\sim\Delta^{-1}$. The green shaded region corresponds to the region where $\mathcal{F}_{\mu\mu}\sim L^{6}$. Hence, as $\Delta\rightarrow 0$, we have $L_{c}\rightarrow\infty$. This corroborates our claim that as $\Delta\rightarrow 0$, we have $\mathcal{F}_{\mu\mu}\sim L^{6}$.
@@
$\zeta(2)=\sum_{n=1}^{\infty}\frac{1}{n^{2}}=\frac{\pi^{2}}{6},$
@@
For very large $L$ and consequently large $m$, we can approximate the sum in bracket of Eq. (28) with Eq. (29). We thus get,
@@
Thus, we have $\mathcal{F}_{\mu\Delta}\sim L^{4}$ for $\Delta\rightarrow 0$ and $\mu\rightarrow 2$. Now, for the case of $\mathcal{G}$, it can be shown that
@@
$\mathcal{F}_{\mu\mu}$ $\approx\frac{\Delta^{2}}{\pi^{6}}L^{6},\quad\mathcal{F}_{\Delta\Delta}\approx\frac{1}{8}L^{2},$
$\mathcal{F}_{\mu\Delta}$ $\approx-\frac{\Delta}{\pi^{4}}L^{4},\quad\mathcal{G}\approx\frac{(\pi^{2}-8)\Delta^{2}}{\pi^{8}}L^{6}.$ (33)
In Fig. 8, we check the validity of Eq. (33) by comparing it with Eq. (23) in the regime of $\Delta\to 0$ and $\mu=2$. Moreover, in Fig. 9 we present the finite-size scaling using Eq. (23) and find that the scaling exponents match with the ones analytically obtained in Eq. (33).
@@
In Fig. 10 we present (a) $\mathcal{F}_{\mu\mu}$ and $(b)\mathcal{F}_{\Delta\Delta}$ for system-size $L=1000$. We observe a similar nature for other system-sizes as well. Fig. 10 quite clearly shows the transition from topological to trivial phase at $\mu=2.0$. We see in Fig. 10(a) that the system achieves its maximum $\mathcal{F}_{\mu\mu}$ at the point $\mu=2.0$. The quantity $\mathcal{F}_{\mu\mu}$ gradually decreases as we move further into the trivial phase. Due to the apparent fluctuations of $\mathcal{F}_{\mu\mu}$ in the topological phase of the finite-sized systems, particularly in smaller systems, numerical calculations were not pushed in the smaller $\Delta$ regime. We, however, performed the scaling analysis in the trivial phase away from the critical line. In Fig. 10(b), we present the component $\mathcal{F}_{\Delta\Delta}$ of QFIM. The figure clearly shows that this quantity increases as we move nearer to $\Delta=0$ line, across which the system the topological invariant undergoes a discrete jump, i.e., a flip in sign of the winding number, given $|\mu|\leq 2$.
Figure 12: Change in the average particle number $\langle N\rangle$ with change in $\Delta$: The average particle number of the probe state Eq. (16) is numerically obtained and is presented with a fixed $\mu=2$ throughout. This a representative case for three system sizes : (a) $L$ = 400 (dot-dashed line), (b) $L$ = 600 (solid line) and (c) $L$ = 800 (dotted line).
Appendix B Average particle number in the ground state
@@
In Fig. 12, we present the average particle number $\langle N\rangle$ of the probe state, i.e., the ground state of the Kitaev model in one dimension for $\mu=2$. As seen from the Hamiltonian of the Kitaev model, the $p$-wave superconductivity term with coupling $\Delta$ makes the system particle-number non-conserving. This means that when $\Delta$ is set to a non-zero value, the ground state wave-function becomes a superposition of various particle-number sector. When $\Delta$ is set to zero, we note two things: the states with fixed momentum become eigenstates of the Hamiltonian and the value of $\mu$ determines the fixed particle-number sector to which the ground state belongs to. We note that at $\mu=0$ limit, half of the single-particle states have negative eigenenergies. The many-body ground state of the model is constructed by filling the negative energy states via fermions. Since there is a negative sign with $\mu$ in the Hamiltonian, more single-particle eigenstates with negative energies becomes available with increasing $\mu$. Thus, occupation number of the ground state increases gradually with increasing $\mu$. On introducing a non-zero $\Delta$, the ground state no longer has a fixed number of particles, rather it is a superposition of many fixed particle-number states. We observe that at very small values of $\Delta$, the ground state has the fully-filled state with almost unit probability because of a small $\Delta$ and a very large $\mu$. As the value of $\Delta$ increases, it becomes energetically more costly to fill up particles than that for smaller $\Delta$. Due to this, the average particle-number decreases with increase of $\Delta$. Hence, physically it turns out that at the gapless multicritical point ($\mu=2$ and $\Delta=0$), the system is in a trivial band-insulating phase, as all available energy states are filled-up by the fermions. A sufficiently small value of the symmetry-breaking term $\Delta$, i.e., $\Delta^{*}$, ensures the system’s transition to a critical $p$-wave superconducting phase. This, in turn, facilitates the quantum-enhanced sensing despite the fact that adiabatic variation of the parameter, $\Delta$, follows the gapless line throughout.